OpenStudy (anonymous):

What is the distance between points (5,-3) and (0,9) in a coordinate system?

OpenStudy (amistre64):

depends on the coordinate system

OpenStudy (amistre64):

if its a standard basis, then you will use the usual distance formula

OpenStudy (amistre64):

subtract, square, add, sqrt

OpenStudy (anonymous):

Point 1 (5,-3) and point 2 (0,9) are point 1 , x = 5 and y = - 3, point 2, x =0, y = 9 x = 5 - 0 = 5, y = -3 - 9 = -12 distance = √((5)^2 + (-12)^2) = 13 The distance = 13

OpenStudy (amistre64):

(5,-3) -(0, 9) ------- 5,-12 25 + 144 = 196 sqrt(196) = 13

OpenStudy (anonymous):

sebasss I am going to marry you !

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

I feel so dumb. Why are you so smart ?

OpenStudy (anonymous):

Only good at math lmao .

OpenStudy (anonymous):

That's what I need. Studying for a GED is hard, bro

OpenStudy (anonymous):

Oh word, Herd it's really hard.